# State Combinatorial Generalization In Decision Making With Conditional Diffusion Models

### v1

@@ 7.1 Single-agent Environment
HighwayEnv (Leurent, 2018) is a self-driving environment where the agent needs to control a vehicle to navigate between traffic controlled by predefined rules. We specifically look at the Roundabout environment with two types of traffic: cars and bicycles (Visualization in Appendix D.7.1).
@@
State in this environment is a composition of four environment vehicles that are either cars or bicycles and the ego agent, which is always a car. Environment observation contains observability, the locations and speed of the ego and surrounding agent, and whether this agent is a car or a bike (Conditioning). During training time, the environment will only generate traffic of all cars or all bicycles with equal probability. During test time, environments will generate a mixture of cars and bicycles (detailed setup in Appendix D.8). Cars and bicycles have different sizes, max speeds, and accelerations, leading to different behavior patterns. This is an instance of generalizing to OOC states with non-overlapping support.
@@ Results
The conditional diffusion model has almost half the number of crashes and higher reward when zero-shot generalizing to states with mixture traffic. Since we train the diffusion model exclusively on successful PPO trajectories, the training state distribution for diffusion is much narrower compared to that of other online methods. This is particularly interesting since it is widely acknowledged that online models have better generalization compared to offline models (Levine et al., 2020).
@@ 7.2 Multi-agent Environment
The unit types in this experiment are Protoss.Stalker, Protoss.Zealot, and Protoss.Colossus, referred to as $a,b,c$ respectively. We evaluate on two OOC scenarios: (1) (Simple: Different but overlapping support): Train the model on randomly generated combinations ($ABC$) of all units and test it where all the units on the team have same type $(AAA)$, (2) (Hard: Non-overlapping support): the opposite scenario, where we train on teams with only one unit type ($AAA$), but during test-time we see any composition of these three units ($ABC$). More information about our setup could be found in Appendix D.9.
Figure 5: Relative improvement % compared to MAPPO on two SMACv2 scenarios: 3v3 and 5v5. Conditional Diffusion show large improvements over the MAPPO baseline, specially in the hard scenario, where we train on teams with the same unit type only but test on random team compositions.
@@ Results
MAPPO performance drastically dropped in the hard OOC scenario by $55.2\%$ for 5v5 and $50\%$ for 3v3. If we substitute one agent generated by MAPPO with conditional diffusion, the success rate can be improved by $16.7\%$ for 3v3 and $23.1\%$ for 5v5 in hard OOC scenario as shown in Figure 5. Detailed success rates are shown in Table 10 and Table 11.
Takeaway 2: Multi-agent RL, viewed from the perspective of a single ego agent, naturally requires combinatorial generalization to collaborate/compete with different agent types. Compositional complexity can be found in a wide range of distinctly different real-world tasks like driving and multiagent decision-making
@@ 7.3 How Do Conditional Diffusion Models Generalize to OOC States?
Figure 6: Rendering of future states predicted by the diffusion model given different conditionings. The grey box is the current state. Blue backgrounds are conditional on all Squares (long attack range) and pink backgrounds are conditioned on all circles (short attack range). Smaller sizes represent less shield and health. More examples shown in Appendix D.9.6.
We can see that conditionings determine the unit type of agent predicted by the diffusion model and also their behavior pattern. Whereas the current state determines other attributes like initial location and health. Different conditionings will lead to different strategies. The circle unit has attack range 1 and the square unit has attack range 6. For units with short attack ranges, the optimal strategy is to approach their enemies before initiating an attack. Conversely, agents with large attack ranges are advised to attack their enemies from a distance to ensure their own safety. Figure 6 shows that if we condition on all circles, the diffusion model thinks players will form a cluster and if condition on all squares, it will predict the players to attack each other from a distance, aligning well with the optimal policy. This demonstrates conditioned diffusion models’ ability to implicitly decompose states to learn underlying compositions and capture multimodality of different unit behavior in the training data. It also demonstrates its ability to perform state stitching to accurately predict the world model.
Takeaway 3: Conditioned diffusion models show significant promise by effectively decomposing and capturing modes of individual base elements and performing state stitching, which helps them to accurately predict the world dynamics and generalize to OOC scenarios.
@@ 8 Ablations
In this section, we ablate over our design choices to (1) show the necessity of using the inductive bias of state latent vector as conditioning, (2) different model architectures to incorporate conditioning information
@@ 8.1 Necessity of Combinatorial Inductive Bias
We compare trajectories generated by the conditioned and unconditioned diffusion models in this section to demonstrate the importance of using combinatorial latent information as conditioning. In Maze2D (Fu et al., 2020), we formulate the navigation problem as a one-step generation process where the diffusion model learns how to generate an entire valid trajectory without rolling out current action and replan. Since there is only one planning step in this process, the generated trajectory can be seen as the “state” in this setting, where unseen trajectories correspond to unseen states instead of time-horizon trajectory stitching. The inductive bias we use is every training trajectory will pass through three waypoints that equally slice the trajectory. In this case, the set of all waypoints forms the base element set and their combination is the latent vector that determines the shape of a generated maze trajectory. During training, we extract three points that equally slice the trajectory and use them as conditioning. During test time, we specify a new combination of three waypoints we want the generated trajectory to pass.
@@
(b) ID (Unconditioned)
(c) OOD (Unconditioned)
(d) cfg=1.3 (Conditioned)
Figure 7: Trajectories generated in Maze2D for large maze. (a) Samples from the training set. (b) Trajectories directly generated by the unconditioned diffusion model given in distribution start and end positions. (c) Trajectories directly generated by the unconditioned diffusion model on unseen start and end positions. (d) Trajectories directly generated by a conditioned diffusion model using 3 waypoints (black dots) as conditioning with classifier-free guidance (cfg) weight 1.3. For results in medium maze please refer to Appendix D.6.1.
We see that the unconditioned diffusion model successfully generated a trajectory if the start and end positions are in the training set (Figure 7(b)) but failed for unseen start and end points (Figure 7(c)). This demonstrates that unconditioned diffusion struggles to approximate unseen distributions. However, if conditioning on an unseen combination of the three waypoints, the conditioned diffusion model can generate unseen trajectories that still satisfy constraints supported by the training dataset (Figure 7(d)), demonstrating the conditioned diffusion model’s ability to generalize to out-of-combination conditioning.
@@ 8.2 Model Architecture: Attention vs Concatenation
Ablation Takeaway: Conditioned diffusion models, provided with information about the new composition of state, can generate better trajectories than unconditioned diffusion models. Also, cross-attention with the condition vector outperforms simply concatenating it with the time-embedding in most cases.
@@ 9 Conclusions
Despite the success of traditional RL models in decision-making tasks, they still struggle to generalize to unseen state inputs. Most existing work focuses on RL generalization under the assumption that generalization to a different probability density function with the same support or allows unseen base elements but also introduces other potentially unrealistic assumptions. However, we take it further and study the problem of generalization to out-of-support states, out of combination in particular, hoping the model can exploit the compositional nature of our world. We showed how this task is challenging for value-based RL and also how conditioned diffusion models can generalize to unsupported samples. We compare the models in different environments with detailed ablation and analysis, demonstrating how each of these classic environments can be formulated as a state combinatorial problem.
However, one limitation of our setup is we model combinatorial generalization in state space as a combination of base elements, which is valid for many real-world applications where complexity stems from exponentially many combinations but does not cover all cases. Oftentimes the distinction between different objects can be blurry (e.g. would a motorbike be a bike). Additionally, the model has difficulty with zero-shot generalization to unseen base objects. Another constraint is efficiency, as planning with diffusion models in stochastic environments requires denoising a trajectory at each planning step, which can be computationally intensive.
Acknowledgment
@@
Ethics Statement
Developing data-driven decision-making models carries the risk of generating inappropriate or harmful actions. This work presents a conditioned model that can be manipulated through carefully forged conditioning, potentially leading to malicious actions.
Reproducibility
Please refer to Section D.3 for detailed model architecture. Pseudocode can be found in Section D.5. For Maze2D, hyperparameters for training conditioned diffusion model can be found in Section D.6.2. For the Roundabout environment, environment parameters (Section D.7.2), dataset detail (Section D.7.3), baseline details (Section D.8.1), hyperparameters for training conditioned diffusion model (Section D.8.1), and models sizes (Section D.8.2) can be found in appendix D.7. For SMACv2 environment, baseline and data collecting policy (Section D.9.1), hyperparameters for training conditioned diffusion model (Section D.9.1), dataset distribution (Section D.9.2), and detailed success rates (Section D.9.3) can be found in appendix (Section D.9). Model runtime and memory can be found in Section D.10.
@@ Appendix A Related Work
Meta RL
Attached below is a review of RL environments for generalization.
@@ A.1 Generalization in RL
Appendix B Proof of Corollary 5.1
Corollary B.1 (Corollary 5.1).
Suppose the states lie along a linear manifold $\mathcal{M}$ in the state space ${\mathbf{S}}$ and the latent space ${\mathbf{Z}}$ is well constructed so that ${\mathbf{Z}}$ is (affine) isomorphic to $\mathcal{M}$. Let ${\bm{s}}$ be a state in the training set with corresponding latent vector ${\bm{z}}$ and ${\bm{s}}^{\prime}$ be an OOC state with corresponding latent vector ${\bm{z}}^{\prime}$. Then a diffusion model $p_{\theta}$ that is well trained on $P_{train}$ can sample ${\bm{s}}^{\prime}$ with non-zero probability.

### v2

@@ 7.1 Single-agent Environment
HighwayEnv (Leurent, 2018) is a self-driving environment where the agent needs to control a vehicle to navigate between traffic controlled by predefined rules. We specifically look at the Roundabout environment with two types of traffic: cars and bicycles (Visualization in Appendix C.6.1).
State in this environment is a composition of four environment vehicles that are either cars or bicycles and the ego agent, which is always a car. Environment observation contains observability, the locations and speed of the ego and surrounding agent, and whether this agent is a car or a bike (Conditioning). During training time, the environment will only generate traffic of all cars or all bicycles with equal probability. During test time, environments will generate a mixture of cars and bicycles (detailed setup in Appendix C.7). Cars and bicycles have different sizes, max speeds, and accelerations, leading to different behavior patterns. This is an instance of generalizing to OOC states with non-overlapping support.
@@ Results
As shown in Figure 4, the conditional diffusion model has almost half the number of crashes and higher reward when zero-shot generalizing to states with mixture traffic. Results of models with different sizes are shown in Table 5 to eliminate the concern of performance gain due to larger model sizes. Since we train the diffusion model exclusively on successful PPO trajectories, the training state distribution for diffusion is much narrower compared to that of other online methods. This is particularly interesting since it is widely acknowledged that online models have better generalization compared to offline models (Levine et al., 2020).
@@ 7.2 Multi-agent Environment
The unit types in this experiment are Protoss.Stalker, Protoss.Zealot, and Protoss.Colossus, referred to as $a,b,c$ respectively. We evaluate on two OOC scenarios: (1) (Simple: Different but overlapping support): Train the model on randomly generated combinations ($ABC$) of all units and test it where all the units on the team have same type $(AAA)$, (2) (Hard: Non-overlapping support): the opposite scenario, where we train on teams with only one unit type ($AAA$), but during test-time we see any composition of these three units ($ABC$). More information about our setup could be found in Appendix C.8.
Figure 5: Relative improvement % compared to MAPPO on two SMACv2 scenarios: 3v3 and 5v5. Conditional Diffusion shows large and consistent improvements over the MAPPO and other BC and RL baselines, especially in the hard scenario, where we train on teams with the same unit type only but test on random team compositions.
@@ Results
As shown in Table 9 and Table 10, MAPPO performance drastically dropped in the hard OOC scenario by $55.2\%$ for 5v5 and $33.3\%$ for 3v3. If we substitute one agent generated by MAPPO with conditional diffusion, the success rate can be improved by $16.7\%$ for 3v3 and $23.1\%$ for 5v5 in hard OOC scenario as shown in Figure 5.
Takeaway 2: Multi-agent RL, viewed from the perspective of a single ego agent, naturally requires combinatorial generalization to collaborate/compete with different agent types.
@@ 7.3 How Do Conditional Diffusion Models Generalize to OOC States?
Figure 6: Rendering of future states predicted by the diffusion model given different conditionings. The grey box is the current state. Blue backgrounds are conditional on all Squares (long attack range) and pink backgrounds are conditioned on all circles (short attack range). Smaller sizes represent less shield and health. More examples shown in Appendix C.8.6.
We can see that player type conditionings determine the unit type of agent predicted by the diffusion model and also their behavior pattern. Whereas the current state determines other attributes like initial location and health. Different player type conditionings will lead to different strategies. The circle unit has attack range 1 and the square unit has attack range 6. For units with short attack ranges, the optimal strategy is to approach their enemies before initiating an attack. Conversely, agents with large attack ranges are advised to attack their enemies from a distance to ensure their own safety. Figure 6 shows that if we condition on all circles, the diffusion model thinks players will form a cluster and if condition on all squares, it will predict the players to attack each other from a distance, aligning well with the optimal policy. This demonstrates conditioned diffusion models’ ability to implicitly decompose states to learn underlying compositions and capture multimodality of different unit behavior in the training data. It also demonstrates its ability to perform state stitching to accurately predict the underlying MDP for an unseen combination.
Takeaway 3: Conditioned diffusion models show significant promise by effectively decomposing and capturing modes of individual base elements and performing state stitching, which helps them to generalize to OOC scenarios.
@@ 8 Ablations
In this section, we ablate over our design choices to (1) show the necessity of using the inductive bias of the latent vector as conditioning, (2) different model architectures to incorporate conditioning information
@@ 8.1 Necessity of Combinatorial Inductive Bias
We compare trajectories generated by the conditioned and unconditioned diffusion models in this section to demonstrate the importance of using the latent vector $\bm{z}$ that contains the combinatorial latent information as conditioning. In Maze2D (Fu et al., 2020), we formulate the navigation problem as a one-step generation process where the diffusion model learns how to generate an entire valid trajectory without rolling out the current action and replanning. Since there is only one planning step in this process, the generated trajectory can be seen as the “state” in this setting, where unseen trajectories correspond to unseen states instead of time-horizon trajectory stitching. The inductive bias we use is every training trajectory will pass through three waypoints that equally slice the trajectory. In this case, the set of all waypoints forms the base element set and their combination is the latent vector that determines the shape of a generated maze trajectory. During training, we extract three points that equally slice the trajectory and use them as conditioning. During test time, we specify a new combination of three waypoints we want the generated trajectory to pass.
@@
(b) ID
(c) OOC
(d) OOC
Figure 7: Trajectories generated in Maze2D for large maze. (a) Samples from the training set. (b) Trajectories generated by the unconditioned diffusion model given in distribution start and end positions. (c) Trajectories generated by the unconditioned diffusion model on unseen start and end positions. (d) Trajectories generated by a conditioned diffusion model using 3 waypoints (black dots) as conditioning. For results in the medium maze please refer to Appendix C.5.1.
We see that the unconditioned diffusion model successfully generated a trajectory if the start and end positions are in the training set (Figure 7(b)), but failed to do so given unseen start and end positions (Figure 7(c)). We compare the trajectories generated by the unconditioned and conditioned diffusion model given unseen start and end positions in Figure 7(c) and 7(d). The trajectory generated by the conditioned diffusion model accurately follows the given OOC waypoints and is drastically different compared to the one without conditioning. This highlights the conditioned diffusion model’s ability to generalize to OOC conditioning and its strong accuracy in adhering to the provided conditioning.
@@ 8.2 Model Architecture: Attention vs Concatenation
Ablation Takeaway: Conditioned diffusion models, provided with information about the new composition of state, are able to generalize to out-of-combination conditioning and accurately adhere to the provided conditioning. Also, cross-attention with the condition vector outperforms simply concatenating it with the time-embedding in most cases.
@@ 9 Conclusions
Despite the success of traditional RL models in decision-making tasks, they still struggle to generalize to unseen state inputs. Most existing work focuses on RL generalization under the assumption that generalization to a different probability density function with the same support. However, we take it further and study the problem of generalization to out-of-support states, out of combination in particular, hoping the model can exploit the compositional nature of our world. We showed how this task is challenging for value-based RL and also how conditioned diffusion models can generalize to unsupported samples. We compare the models in different environments with detailed ablation and analysis, demonstrating how each of these classic environments can be formulated as a state combinatorial problem.
However, one limitation of our setup is we model combinatorial generalization in state space as a combination of base elements, which is valid for many real-world applications but does not cover all cases. Additionally, the model has difficulty with zero-shot generalization to unseen base objects. Another constraint is efficiency, as planning with diffusion models in stochastic environments requires denoising a trajectory at each planning step, which can be computationally intensive. Various approaches have been proposed to accelerate the sampling process in diffusion models via advanced ODE solvers or knowledge distillation techniques (Song et al., 2020a; Karras et al., 2022; Kim et al., 2023; Song et al., 2023). Incorporating these works into diffusion planning would be a promising future direction to address rollout inefficiency.
Broader Impact Statement
Developing data-driven decision-making models carries the risk of generating inappropriate or harmful actions. This work presents a conditioned model that can be manipulated through carefully forged conditioning, potentially leading to malicious actions. Our experiments were conducted in simulated environments, and testing this model in the real world without safety restrictions could be dangerous.
Acknowledgments
@@ Appendix A Related Work
We introduced related problem setup and generalization methods in Section 2. Attached below is a review of RL environments for generalization.
@@ A.1 Generalization in RL
Appendix B Proof of Theorem 5.1
In this section, we provide the full statement and proofs of Theorem 5.1. First, we establish the following assumptions.
Practically, when the element set is discrete, an embedding $\phi:{\mathbf{E}}^{n}\rightarrow\mathbb{R}^{m}$ is applied to map a discrete representation of latent elements to a continuous embedding space.
Assumption B.1 (Linear manifold assumption).
Assume the states lie along a linear manifold $\mathcal{M}$ in the ambient state space ${\mathbf{S}}\subset\mathbb{R}^{d}$ and the encoded latent space $\phi({\mathbf{Z}})\subset\mathbb{R}^{m}$ with $m<d$ is well constructed so that $\phi({\mathbf{Z}})$ is (affine) isomorphic to $\mathcal{M}$.
Here we can interpret the linear manifold $\mathcal{M}$} as the subspace containing all “in-combination” (i.e. in support or out-of-combination) data points. Recall we can parametrize the DDPM learned reverse process of a diffusion model as
$p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t}):=\mathcal{N}({\bm{x}}_{t-1};\bm{\mu_{\theta}}({\bm{x}}_{t},t),\sigma_{t}I)$ (2)
where $\sigma_{t}$ is a pre-fixed noise schedule. Given Assumption B.1, we denote the projection operation onto the manifold $\mathcal{M}$} as $W_{\mathcal{M}}$ and the projector onto its orthogonal complement $\mathcal{M}$$^{\perp}$ as $W_{\perp}$. We then further assume the diffusion model is well-trained, in particular, it should satisfy the following assumptions:
Assumption B.2 (Well-trained diffusion model).
We assume access to a well-trained DDPM diffusion model that satisfies:
  1. 1.
    Block-wise bi-Lipschitz with contraction: For all ${\bm{x}},{\bm{y}}\in{\mathbf{S}}$ and all timesteps $t$, there exists constants
    $0<A_{t}^{\mathcal{M}}\leq B_{t}^{\mathcal{M}}<\infty,\qquad 0<A_{t}^{\perp}\leq B_{t}^{\perp}<1$ (3)
    such that
    $A_{t}^{\mathcal{M}}\|W_{\mathcal{M}}({\bm{x}}-{\bm{y}})\|$ $\leq\|W_{\mathcal{M}}(\bm{\mu_{\theta}}({\bm{x}},t)-\bm{\mu_{\theta}}({\bm{y}},t))\|\leq B_{t}^{\mathcal{M}}\|W_{\mathcal{M}}({\bm{x}}-{\bm{y}})\|$ (4)
    $A_{t}^{\perp}\|W_{\perp}({\bm{x}}-{\bm{y}})\|$ $\leq\|W_{\perp}(\bm{\mu_{\theta}}({\bm{x}},t)-\bm{\mu_{\theta}}({\bm{y}},t))\|\leq B_{t}^{\perp}\|W_{\perp}({\bm{x}}-{\bm{y}})\|$
  2. 2.
    Block preserving denoising: For all timesteps $t$, $\bm{\mu_{\theta}}(\cdot,t)$ preserves the block decomposition. I.e.
    $W_{\mathcal{M}}\bm{\mu_{\theta}}(W_{\perp}{\bm{x}},t)=W_{\perp}\bm{\mu_{\theta}}(W_{\mathcal{M}}{\bm{x}},t)=\mathbf{0}$ (5)
    for all $x\in{\mathbf{S}}$.
The first assumption suggests that a well-trained diffusion model will contract the components that are orthogonal to the manifold and pull the samples towards the manifold (i.e. towards “in-combination” samples). The second assumption indicates that the learned reverse mean should not mix the two orthogonal components.
Based on these two assumptions, we can now prove the following lemmas that will be helpful for our proof to Theorem 5.1.
Lemma B.3.
If $X\in\mathbb{R}^{d}$ has density lower bound
$p_{X}({\bm{x}})\geq c(2\pi)^{-d/2}(\sigma^{\mathcal{M}})^{-m}(\sigma^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{x}}-\mu)\|^{2}}{2(\sigma^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{x}}-\mu)\|^{2}}{2(\sigma^{\perp})^{2}}\right)}$ (6)
Let $g$ be a function that satisfies block-wise bi-Lipschitz with contraction and block reservation from Assumption B.2 with Lipschitz constants $A^{\mathcal{M}},B^{\mathcal{M}},A^{\perp},B^{\perp}$ and $Y=g(X)$, then for constant $c$ we can have
$p_{Y}({\bm{y}})\geq c(2\pi)^{-d/2}(\sigma^{\mathcal{M}}B^{\mathcal{M}})^{-m}(\sigma^{\perp}B^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\mathcal{M}}A^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\perp}A^{\perp})^{2}}\right)}$ (7)